7. Find the quadriceps muscle force (F3) needed to hold the leg in the horizontal position during an isometric leg extension exercise, given that F2 = 180 N,F1-75 N and θ angle of pull of the quads with respect to the horizontal). 25° (ie, the F, 10am 20
Answer #1


In equilibrium net torque = 0


Fy3*r3 - F1*r1 - F2*r2 = 0


r1= 60 cm = 0.6 m


r2 = 20 cm = 0.2 m

r3 = 10 cm = 0.1 m


Fy3*0.1 - 75*0.2 - 180*0.6 = 0


Fy3 = 1230 N


Fy3 = F3*sintheta

F3 = Fy3/sintheta

F3 = 1230/sin25

F3 = 2910 N


-===================